# QMGeo: Differentially Private Federated Learning via Stochastic Quantization with Mixed Truncated Geometric Distribution

### v1

@@ I-A Related Work
There exists studies that jointly consider reducing communication cost by quantization and providing DP guarantees. For instance, the authors in [12] propose applying the binomial mechanism with quantization to achieve both communication efficiency and differential privacy, with an improved analysis showing better utility for the binomial mechanism. [13] efficiently discretizes and flattens the model updates from the users and adds discrete Gaussian noise. However, these mechanisms still rely heavily on additional noise injection to achieve differential privacy.
@@ II Federated Learning Protocol
$\boldsymbol{w}^{t+1}=\boldsymbol{w}^{t}-\eta\sum_{n=1}^{N}\Delta\boldsymbol{w}_{n}(t).$ (4)
The updated global model parameter $w^{t+1}$ is then broadcast to all the users before the next iteration.
@@ III QMGeo: Quantization Method with Mixed Geometric Distribution
$Q_{\text{MGeo}}(w)$ takes an integer $w$ as input and outputs a value from the set of quantization levels $\left\{Bin(0),Bin(1),\dots,Bin(r),\dots,Bin(R-1)\right\}$ as the quantized output. The output value, i.e., the chosen quantization level $Bin(r)$, is sampled from a mixed truncated geometric distribution determined by the value of the input integer $w$.
@@ III-B Mixed Truncated Geometric Distribution
$X_{\text{MTGeo}}=\left\{\begin{array}[]{rcl}Bin\left(r-(X_{{Geo1}}-1)\right)&\text{w.p.}&p_{\text{mix}}\\ Bin\left(r+X_{{Geo2}}\right)&\text{otherwise}\\ \end{array}\right.\$ (9)
where $p_{\text{mix}}=\frac{w-Bin(r)}{Bin(r+1)-Bin(r)}$, and $X_{{TGeo1}}-1$ and $X_{{TGeo2}}-1$ are random variables sampled from two different truncated geometric distributions, ${Geo1}(p)$ and ${Geo2}(p)$. The mixture probability is determined by the distance from $w$ to $Bin(r)$ and $Bin(r+1)$ respectively, where $Bin(r)$ and $Bin(r+1)$ are the neighboring quantization levels to $w$.
Since both the input and the output of the quantization scheme has a range of $[-W_{\text{max}},W_{\text{max}}]$, we are forced to use truncated geometric distributions to sample to decided quantization level. Thus, $X\sim{TGeo1}(p)$ is supported only on $\{1,2,\dots,r+1\}$ and $X\sim\text{TGeo2}(p)$ is supported only on $\{1,2,\dots,R-r-1\}$. We obtain the truncated geometric distributions by normalizing the probability mass on the support back to $1$, where the geometric distribution itself is parameterized by the success probability $p$:
$Pr\{X_{{Geo1}}=k\}=p(1-p)^{k-1}\cdot\frac{1}{1-(1-p)^{r}},$ (10)
$Pr\{X_{{Geo2}}=k\}=p(1-p)^{k-1}\cdot\frac{1}{1-(1-p)^{R-r-2}}.$ (11)
@@
Figure 1: How $\mathcal{Q}_{\text{MGeo}}(\cdot)$ is demonstrated in this figure. The dotted line shows the probability of the input value $w$ being quantized to the corresponding quantization level. As shown in the figure, every quantization level is assigned with a non-zero probability. The larger $p$ is, the more skewed the distribution becomes.
@@ IV-A Differential privacy of the stochastic $k$-level quantization
For number of quantization levels larger than 3, the stochastic $k$-level quantization faces a difficulty providing DP, as it could only quantize values to their neighboring quantization levels. When analysing the DP performance, it is obvious that if $g$ and $g^{{}^{\prime}}$ do not lie in neighboring intervals, there will always be a choice of $v$ that achieves the worst case rendering one of the probability terms to 0 and nonzero for the other. For instance, when $v$ is a neighboring quantization level $v=Bin(r)$ to $g$ but not for $g^{{}^{\prime}}$, $\epsilon$ cannot be unbounded. An example is provided in Fig 2, where the intuition is that, since the quantization method quantizes to only neighboring quantization levels, there is not enough randomness to achieve DP for a number of quantization levels larger than 3. However, for ternary quantization, such method could provide DP guarantees to some extent [15]:
$e^{\epsilon}Pr[\mathcal{Q}(g^{{}^{\prime}})=v]$ $\leq Pr[\mathcal{Q}(g)=v]$ (17)
$e^{\epsilon}\times 0$ $\leq\frac{g-v}{Bin(r+1)-v}.$ (18)
Figure 2: This figure illustrates an example of the stochastic $k$-level quantization failing to provide DP. The solid orange and blue lines are the possible output quantization levels for $g$ and $g^{{}^{\prime}}$ respectively. It is obvious that, by observing the output of the quantization, any adversary could distinguish betwen $g$ and $g^{{}^{\prime}}$. Any grey dotted lines are quantization levels that lies neither in the range of $Q(g)$ or $Q(g^{{}^{\prime}})$, thus each of them being chosen as $v$ would render the $\epsilon$ unbounded.
@@ IV-B1 $\epsilon$-DP
(19) should hold for any $v$, where $g$ and $g^{{}^{\prime}}$ are entries of gradients acquired from arbitrary neighboring datasets of a particular user. We note that the worst case scenario occurs when $v$ is either $B(0)$ or $B(R-1)$, $g$ and $g^{{}^{\prime}}$ is equal to $B(0)$ and $B(R-1)$ respectively. Without loss of generality, we assume $v=B(0)$, $g=B(0)$ and $g^{{}^{\prime}}=B(R-1)$. Substituting the values in (19) with the probability mass of the mixed truncated geometric distribution, we obtain
$\frac{Pr[Q_{\text{MGeo}}(g)=v]}{Pr[Q_{\text{MGeo}}(g^{{}^{\prime}})=v]}$ $\leq e^{\epsilon}$ (20)
$\frac{\frac{1}{2}}{\frac{1}{2}p(1-p)^{R-2}}$ $\leq e^{\epsilon}$ (21)
$\epsilon$ $\leq-\ln{p(1-p)^{R-2}}$
$=-(\ln{p}+(R-2)\ln{(1-p))}.$
We demonstrate the trade-off between the achieved DP level $\epsilon$ and the parameters of the $QMGeo(\cdot)$, the number of quantization levels $r$ and the success probability $p$ of the mixed truncated geometric distribution. Fig 3 demonstrates how the achieved DP level $\epsilon$ changes as the parameter $p$ of $Q_{\text{QMGeo}}(\cdot)$ changes for fixed numbers of quantization levels. $\epsilon$ increases rapidly as $p$ approaches 1, where at $p=1$, the $Q_{\text{QMGeo}}(\cdot)$ reduces to a conventional stochastic $r$-level quantization which could not achieve DP except for when quantizing to 3 levels or less. As shown in (20), $\epsilon$ increases linearly with the number of quantization levels $R$.
Figure 3: The curves displayed in this figure each corresponds to a different number of quantization levels. The y-axis shows the achieved $\epsilon$-DP level, while the x-axis shows the parameter $p$ of $Q_{\text{QMGeo}}(\cdot)$, the success rate of the mixed truncated geometric distribution.
@@ IV-B2 $(\alpha,\epsilon)$-RDP
With the quasi-convex property of the Rényi divergence, we can show the worst case scenario is attained by the extremal points where two input elements each sit at $Bin(0)$ and $Bin(R-1)$, where the two discrete distributions are essentially a standard truncated geometric distribution $X$ and a slightly altered one, $X^{{}^{\prime}}=-X+R+1$, sharing the same parameter of success rate $p$ and supported on the same set of quantization levels. We then substitute parameters $R$, $p$ from the $\mathcal{Q}_{\text{MGeo}(\cdot)}$ into (13) and obtain
@@
$=\frac{1}{\alpha-1}\log\left\{\left(\sum_{i=1}^{R}\frac{\left(pq^{i-1}\frac{1}{1-q^{R-1}}\right)^{\alpha}}{\left(q^{R+1-i}p\frac{1}{1-q^{R-1}}\right)^{\alpha-1}}\right)\right\}$ (23)
$=\frac{1}{1-\alpha}\log\left\{\left(\frac{pq^{-2\alpha+(1-\alpha)R+1}}{1-q^{R-1}}\right)\frac{\alpha\left(q^{(2\alpha-1)R}-1\right)}{q^{(2\alpha-1)}-1}\right\}$ (24)
@@
$\epsilon(\alpha)\leq\frac{1}{1-\alpha}\log\left\{\frac{pq^{-2\alpha+(1-\alpha)R+1}}{1-q^{R-1}}\cdot\frac{\alpha\left(q^{(2\alpha-1)R}-1\right)}{q^{(2\alpha-1)}-1}\right\}.$ (25)
Figure 4: This figure show cases how $\epsilon(\alpha)$ changes as $\alpha$ changes. The y-axis is the achieved $\epsilon$, the x-axis is the corresponding $\alpha$. The curve is obtained with a $\mathcal{Q}_{\text{MGeo}}(\cdot)$ with $r=8$ $p=0.5$.
Fig. 4 plots the $\epsilon(\alpha)$ curve obtained with a $\mathcal{Q}_{\text{MGeo}}(\cdot)$ with $R=8$ and $p=0.5$.
@@ IV-C2 $(\alpha,\epsilon)$-RDP
$\epsilon(\alpha)\leq$ $\kappa^{2}d\frac{1}{1-\alpha}\times$ (27)
$\log\left\{\frac{pq^{-2\alpha+(1-\alpha)R+1}}{1-q^{R-1}}\cdot\frac{\alpha\left(q^{(2\alpha-1)R}-1\right)}{q^{(2\alpha-1)}-1}\right\}.$
@@ V Optimality Gap
We next characterize how the perturbation introduced by the randomized quantization method impact the convergence performance by characterizing the optimality gap between $F(\boldsymbol{w}_{T})$ (the loss value achieved at global iteration $T$) and $F^{*}$, the optimal loss value.
@@ VI Numerical Results
The following results are obtained on MNIST dataset (10 class handwritten digit dataset) with a total of 60000 samples. The dataset is distributed as follows: 90% of the data are randomly split into 5 users, 10% of the data are reserved at the PS for evaluation. The FL framework trains a multi-layer perceptron with one hidden layer of 32 nodes. We first perform principal component analysis on the dataset and reduce the dimensionality of the input data to 100 to speed up the computation. The dimension of the model update vector is $d=3562$.
@@
We use (27) to calculate the $\epsilon(\alpha)$ per round achieved by the $Q_{\text{QMGeo}}(\cdot)$ method. Fig 5 demonstrates the accuracy comparison between setting $p=0.9$ and $p=0.5$, when the number of quantization levels is fixed at $R=8$. The dotted line is a base case in which no quantization is performed. All curves approach a similar accuracy level, indicating that quantization does not result in performance degradation in terms of accuracy at least in the considered setup. In terms of privacy guarantees, using $Q_{\text{QMGeo}}(\cdot)$ with $p=0.9$ and $R=8$ achieved $\epsilon=2.626$ per round, with $p=0.5$ and $R=8$ achieved $\epsilon=0.784$ per round.
Fig 6 demonstrates the accuracy comparison between setting $R=16$ and $R=8$, while fixing the parameter $p=0.9$. Using $Q_{\text{QMGeo}}(\cdot)$ with $p=0.9$ and $R=16$ achieved $\epsilon=4.492$ per round, with $p=0.9$ and $R=8$ achieved $\epsilon=2.626$ per round.
Figure 5: The two solid line curves displayed in this figure corresponds to a $p=0.5$ and $p=0.9$ in $Q_{\text{QMGeo}}(\cdot)$ respectively, with the number of quantization levels $R=8$. The dotted curve is the baseline curve where we do not apply any quantization. The y-axis shows the accuracy obtained at the global model using the hold out test set, while the x-axis shows the number of communication rounds. The $\epsilon$ labeled in the legend is acquired using (27), fixing $\alpha=2$.
Figure 6: The two curves displayed in this figure corresponds to a $R=16$ and $R=8$ in $Q_{\text{QMGeo}}(\cdot)$ respectively, with the parameter of success rate $p=0.9$. The y-axis shows the accuracy obtained at the global model using the hold out test set, while the x-axis shows the number of communication rounds. The $\epsilon$ labeled in the legend is acquired using (27), fixing $\alpha=2$.

### v2

@@ I-A Related Work
There exist studies that jointly consider reducing communication cost by quantization and providing DP guarantees. For instance, the authors in [12] propose applying the binomial mechanism with quantization to achieve both communication efficiency and differential privacy, with an improved analysis showing better utility for the binomial mechanism. [13] efficiently discretizes and flattens the model updates from the users and adds discrete Gaussian noise. However, these mechanisms still rely heavily on additional noise injection to achieve differential privacy.
@@ II Federated Learning Protocol
$\boldsymbol{w}_{t+1}=\boldsymbol{w}_{t}-\eta\sum_{n=1}^{N}\Delta\boldsymbol{w}_{n}(t).$ (4)
The updated global model parameter $\boldsymbol{w}_{t+1}$ is then broadcasted to all the users before the next iteration.
@@ III QMGeo: Quantization Method with Mixed Geometric Distribution
$Q_{\text{MGeo}}(w)$ takes a scalar $w$ as input and outputs a value from the set of quantization levels $\left\{Bin(0),Bin(1),\dots,Bin(r),\dots,Bin(R-1)\right\}$ as the quantized output. The output value, i.e., the chosen quantization level $Bin(r)$, is sampled from a mixed truncated geometric distribution determined by the value of the input scalar $w$.
@@ III-B Mixed Truncated Geometric Distribution
$X_{\text{MTGeo}}=\left\{\begin{array}[]{rcl}Bin\left(r-(X_{{TGeo1}}-1)\right)&\text{w.p.}&p_{\text{mix}}\\ Bin\left(r+X_{{TGeo2}}\right)&\text{otherwise}\\ \end{array}\right.\$ (9)
where $p_{\text{mix}}=\frac{Bin(r+1)-w}{Bin(r+1)-Bin(r)}$, and $X_{{TGeo1}}$ and $X_{{TGeo2}}$ are random variables sampled from two different truncated geometric distributions, ${TGeo1}(p)$ and ${TGeo2}(p)$. The mixture probability is determined by the distance from $w$ to $Bin(r)$ and $Bin(r+1)$ respectively, where $Bin(r)$ and $Bin(r+1)$ are the neighboring quantization levels to $w$.
Since both the input and the output of the quantization scheme has a range of $[-W_{\text{max}},W_{\text{max}}]$, we are forced to use truncated geometric distributions to sample the decided quantization level. Thus, $X\sim{TGeo1}(p)$ is supported only on $\{1,2,\dots,r+1\}$ and $X\sim{TGeo2}(p)$ is supported only on $\{1,2,\dots,R-r-1\}$. We obtain the truncated geometric distributions by normalizing the probability mass on the support back to $1$, where the geometric distribution itself is parameterized by the success probability $p$:
$Pr\{X_{{TGeo1}}=k\}=p(1-p)^{k-1}\cdot\frac{1}{1-(1-p)^{r+1}},$ (10)
$Pr\{X_{{TGeo2}}=k\}=p(1-p)^{k-1}\cdot\frac{1}{1-(1-p)^{R-r-1}}.$ (11)
@@
Figure 1: How $\mathcal{Q}_{\text{MGeo}}(\cdot)$ quantizes a scalar input value $w$ is demonstrated in this figure. The dotted line shows the probability of the input value $w$ being quantized to the corresponding quantization level. As shown in the figure, each quantization level is assigned with a non-zero probability. The larger $p$ is, the more skewed the distribution becomes.
@@ IV-A Differential privacy of the stochastic $k$-level quantization
For number of quantization levels larger than 3, the stochastic $k$-level quantization faces a difficulty providing DP, as it could only quantize values to their neighboring quantization levels. When analysing the DP performance, it is obvious that if $g$ and $g^{{}^{\prime}}$ do not lie in neighboring intervals, there will always be a choice of $v$ that achieves the worst case rendering one of the probability terms to 0 and nonzero for the other. For instance, when $v$ is a neighboring quantization level $v=Bin(r)$ to $g$ but not for $g^{{}^{\prime}}$, $g$ will be quantized to $v$ with a non-zero probability $p_{mix}$, while $g^{{}^{\prime}}$ will never be quantized to $v$. In this case, $\epsilon$ cannot be bounded. An example is provided in Fig 2, where the intuition is that, since the quantization method quantizes to only neighboring quantization levels, there is not enough randomness to achieve DP for a number of quantization levels larger than 3:
$e^{\epsilon}Pr[\mathcal{Q}(g^{{}^{\prime}})=v]$ $\geq Pr[\mathcal{Q}(g)=v]$ (17)
$e^{\epsilon}\times 0$ $\geq p_{mix}.$ (18)
However, for ternary quantization, such method could provide DP guarantees to some extent [15].
Figure 2: This figure illustrates an example of the stochastic $k$-level quantization failing to provide DP. The solid orange and blue lines are the possible output quantization levels for $g$ and $g^{{}^{\prime}}$ respectively. It is obvious that, by observing the output of the quantization, any adversary could distinguish betwen $g$ and $g^{{}^{\prime}}$. Any grey dotted lines are quantization levels that lies neither in the range of $Q(g)$ or $Q(g^{{}^{\prime}})$, thus any of them being chosen as $v$ would render the $\epsilon$ unbounded.
@@ IV-B1 $\epsilon$-DP
(19) should hold for any $v$, where $g$ and $g^{{}^{\prime}}$ are entries of gradients acquired from arbitrary neighboring datasets of a particular user. We note that the worst case scenario occurs when $v$ is either $Bin(0)$ or $Bin(R-1)$, $g$ and $g^{{}^{\prime}}$ are equal to $Bin(0)$ and $Bin(R-1)$ respectively. Without loss of generality, we assume $v=Bin(0)$, $g=Bin(0)$ and $g^{{}^{\prime}}=Bin(R-1)$. Substituting the values in (19) with the probability mass of the mixed truncated geometric distribution, we obtain
$e^{\epsilon}$ $\geq\frac{Pr[Q_{\text{MGeo}}(g)=v]}{Pr[Q_{\text{MGeo}}(g^{{}^{\prime}})=v]}$ (20)
$e^{\epsilon}$ $\geq\frac{\frac{1}{2}}{\frac{1}{2}p(1-p)^{R-2}\frac{1}{1-(1-p)^{R-1}}}$ (21)
$\epsilon$ $\geq-\ln{p(1-p)^{R-2}}+\ln{(1-(1-p)^{R-1})}$
$=-(\ln{p}+(R-2)\ln{(1-p))}+\ln{(1-(1-p)^{R-1})}.$
We demonstrate the trade-off between the achieved DP level $\epsilon$ and the parameters of the $QMGeo(\cdot)$ i.e., the number of quantization levels $r$ and the success probability $p$ of the mixed truncated geometric distribution. Fig 3 demonstrates how the achieved DP level $\epsilon$ changes as the parameter $p$ of $Q_{\text{QMGeo}}(\cdot)$ changes for fixed numbers of quantization levels. $\epsilon$ increases rapidly as $p$ approaches 1, where at $p=1$, the $Q_{\text{QMGeo}}(\cdot)$ reduces to a conventional stochastic $r$-level quantization which could not achieve DP except for when quantizing to 3 levels or less. As shown in (20), $\epsilon$ increases linearly with the number of quantization levels $R$.
Figure 3: $\epsilon$-DP as a function of the parameter $p$ of $Q_{\text{QMGeo}}(\cdot)$, the success rate of the mixed truncated geometric distribution. Each curve displayed in this figure corresponds to a different number of quantization levels.
@@ IV-B2 $(\alpha,\epsilon)$-RDP
With the quasi-convex property of the Rényi divergence, we can show that the worst case scenario is attained by the extremal points where two input elements each sit at $Bin(0)$ and $Bin(R-1)$, where the two discrete distributions are essentially a standard truncated geometric distribution $X$ and a slightly altered one, $X^{{}^{\prime}}=-X+R+1$, sharing the same parameter of success rate $p$ and supported on the same set of quantization levels. We then substitute parameters $R$, $p$ from the $\mathcal{Q}_{\text{MGeo}(\cdot)}$ into (13) and obtain
@@
$=\frac{1}{\alpha-1}\log\left\{\frac{1}{2}\frac{(1-q^{R-1})^{\alpha-1}}{(q^{R}p)^{\alpha-1}}+\frac{1}{2}(pq^{R-1}\frac{1}{1-q^{R-1}})^{\alpha}\right.$ (23)
$+\left.\frac{pq^{-2\alpha+(1-\alpha)R+1}}{2(1-q^{R-1})}\cdot\frac{q^{4\alpha-2}(1-q^{(2\alpha-1)(R-2)})}{1-\ q^{4\alpha-2}}\right\}$
@@
$\epsilon(\alpha)$ $\geq\frac{1}{\alpha-1}\log\left\{\frac{1}{2}\frac{(1-q^{R-1})^{\alpha-1}}{(q^{R}p)^{\alpha-1}}+\frac{1}{2}(pq^{R-1}\frac{1}{1-q^{R-1}})^{\alpha}\right.$ (24)
$+\left.\frac{pq^{-2\alpha+(1-\alpha)R+1}}{2(1-q^{R-1})}\cdot\frac{q^{4\alpha-2}(1-q^{(2\alpha-1)(R-2)})}{1-\ q^{4\alpha-2}}\right\}.$
Figure 4: Plot of $\epsilon(\alpha)$ vs. $\alpha$. The y-axis is the achieved $\epsilon$, and the x-axis is the corresponding $\alpha$. The curve is obtained with a $\mathcal{Q}_{\text{MGeo}}(\cdot)$ with $R=8$ $p=0.5$.
Fig 4 plots the $\epsilon(\alpha)$ curve obtained with a $\mathcal{Q}_{\text{MGeo}}(\cdot)$ with $R=8$ and $p=0.5$.
@@ IV-C2 $(\alpha,\epsilon)$-RDP
$\epsilon(\alpha)\geq$ $\kappa^{2}d\frac{1}{\alpha-1}\times$ (26)
$\log\left\{\frac{1}{2}\frac{(1-q^{R-1})^{\alpha-1}}{(q^{R}p)^{\alpha-1}}+\frac{1}{2}(pq^{R-1}\frac{1}{1-q^{R-1}})^{\alpha}\right.$
$+\left.\frac{pq^{-2\alpha+(1-\alpha)R+1}}{2(1-q^{R-1})}\cdot\frac{q^{4\alpha-2}(1-q^{(2\alpha-1)(R-2)})}{1-\ q^{4\alpha-2}}\right\}.$
@@ V Optimality Gap
We next characterize how the perturbation introduced by the randomized quantization method impacts the convergence performance by characterizing the optimality gap between $F(\boldsymbol{w}_{T})$ (the loss value achieved at global iteration $T$) and $F^{*}$, the optimal loss value.
@@ VI Numerical Results
The following results are obtained on MNIST dataset (10 class handwritten digit dataset) with a total of 60000 samples. The dataset is distributed as follows: 90% of the data are randomly split into 5 users, and 10% of the data are reserved at the PS for evaluation. The FL framework trains a multi-layer perceptron with one hidden layer of 32 nodes. We first perform principal component analysis on the dataset and reduce the dimensionality of the input data to 100 to speed up the computation. The dimension of the model update vector is $d=3562$.
@@
We use (26) to calculate the $\epsilon(\alpha)$ per round achieved by the $Q_{\text{QMGeo}}(\cdot)$ method. Fig 5 demonstrates the accuracy comparison between setting $p=0.9$ and $p=0.5$, when the number of quantization levels is fixed at $R=8$. The dotted line is a baseline case in which no quantization is performed. All curves approach a similar accuracy level, indicating that quantization does not result in performance degradation in terms of accuracy at least in the considered setup. In terms of privacy guarantees, using $Q_{\text{QMGeo}}(\cdot)$ with $p=0.9$ and $R=8$ achieves $\epsilon=1.807$ per round, and $Q_{\text{QMGeo}}(\cdot)$ with $p=0.5$ and $R=8$ achieves $\epsilon=0.564$ per round.
Fig 6 demonstrates the accuracy comparison between setting $R=16$ and $R=8$ for fixed parameter $p=0.9$. Using $Q_{\text{QMGeo}}(\cdot)$ with $p=0.9$ and $R=16$ achieves $\epsilon=3.673$ per round, while $Q_{\text{QMGeo}}(\cdot)$ with $p=0.9$ and $R=8$ leads to $\epsilon=1.807$ per round.
Figure 5: The two solid line curves displayed in this figure correspond to $p=0.5$ and $p=0.9$ in $Q_{\text{QMGeo}}(\cdot)$ respectively, with the number of quantization levels $R=8$. The dotted curve is the baseline curve where we do not apply any quantization. The y-axis shows the accuracy obtained at the global model using the hold-out test set, while the x-axis shows the number of communication rounds. The $\epsilon$ labeled in the legend is acquired using (26), fixing $\alpha=2$.
Figure 6: The two curves displayed in this figure correspond to $R=16$ and $R=8$ in $Q_{\text{QMGeo}}(\cdot)$ respectively, with the parameter of success rate $p=0.9$. The y-axis shows the accuracy obtained at the global model using the hold-out test set, while the x-axis shows the number of communication rounds. The $\epsilon$ labeled in the legend is acquired using (26), fixing $\alpha=2$.